The table to the right shows the total runs scored by 16 baseball teams during one year.

a. Determine the percentile rank for the runs scored by team

Upper NN.

b. Determine the percentile rank for the runs scored by team

Upper EE.

c. Determine the percentile rank for the runs scored by team

Upper BB.

Team

Runs

Team

Runs

A

761761

I

664664

B

760760

J

645645

C

755755

K

627627

D

752752

L

617617

E

750750

M

615615

F

715715

N

600600

G

689689

O

594594

H

671671

P

580580
Answer #1

percentile ranking is a method of ranking that uses scale of 100.

percentile is calculated as follows

(Number of teams behind / total number of teams)*100

Now here total number of baseball team is 16 and as we can see all the teams are in natural ranking.

a) Here we have to determine percentile rank of team N.

Now the rank of team N is 14 .

hence, the number of team behind it is (16-14)=2

so percentile rank =(2/16)*100 = 12.5

b) Her we have to determine percentile rank for team E.

the rank of team E is 5.

so number of team behind it is (16-5)=11

Hence percentile rank for team E is (11/16)*100 = 68.75

c)Now percentile rank scored by team B is

rank of team B is 2

Number of team behind it =(16-2)=14

hence percentile rank for team B =(14/16)*100=87.5